OpenStudy (anonymous):

Hi my daughter is having trouble with an assignment. She is asked to write in her journal how to find a common denominator? She has to construct cubes and the first problem is 1/3 and 2/12

OpenStudy (anonymous):

ok so to find a common denominator first you need to know what a denominator is. a denominator is the bottom number under a fraction

OpenStudy (anonymous):

so all you have to do is find the first common times that 3 and 12 go into

OpenStudy (anonymous):

3 goes into 12 pretty easy when you times it by 4 so whatever you do to the bottom you do to the top. so it would be 4/12 and 2/12 because we didn't have to times the 12 by anything it stayes the same

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

ARE you there?

OpenStudy (anonymous):

understand. But she is not allowed to use LCM on her assignment. The problem I wrote are two different problems. Can you still help us?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

LCM? sorry i havn't heard that term or i musv't forgot? sorry

OpenStudy (anonymous):

Least common denominator is the abbreviation.

OpenStudy (anonymous):

oh ok but isn't that what is asked?

OpenStudy (anonymous):

I'm haven't done fractions in so long sorry. She is a fourth grader and I'm a little rusty.

OpenStudy (anonymous):

You need the LCD, not the LCM right?

OpenStudy (anonymous):

Yeah sorry

OpenStudy (anonymous):

And what is the problem again?

OpenStudy (anonymous):

She has to find the common denominator and use blocks to shade in the fraction. 1/3 is the first and 2/12 is second

OpenStudy (anonymous):

common denominator of 1/3 and 2/12?

OpenStudy (anonymous):

The least common denominator of 1/3 and 2/12 is 6.

OpenStudy (anonymous):

i CAN'T DO ANYTHING!!! :'(

OpenStudy (anonymous):

dont listen to arada he is always wrong he knows nothing about math

OpenStudy (anonymous):

\[\frac{2}{12} = \frac{2*1}{2*6} = \frac{1}{6}\] \[\frac{1}{3} = \frac{2}{2}*\frac{1}{3} = \frac{2}{6}\]

OpenStudy (anonymous):

Okay does she shade in the 6 blocks and leave one blank?

OpenStudy (anonymous):

Do you have a digital camera? If so can you take a small picture of the problem and attach it here? I'm not sure what is being asked..

OpenStudy (anonymous):

'Your daughter'? Your lies don't fall me! This is your homework, isn't it!

OpenStudy (anonymous):

fool* fail

OpenStudy (anonymous):

Really it's her homework. She has a journal book and she must explain her answers.

OpenStudy (anonymous):

So I still don't understand the question, what is she asked to do?

OpenStudy (anonymous):

She has to explain by using model blocks or cubes and show her work

OpenStudy (anonymous):

I'm not sure how you'd do that.

OpenStudy (anonymous):

I'm trying to understand what the teacher is asking. Not very clear about this assignment. What do I do?

OpenStudy (anonymous):

As I said, can you attach a picture of the problem?

OpenStudy (anonymous):

Otherwise ask her teacher about it.

OpenStudy (anonymous):

How do I attach the problem?

OpenStudy (anonymous):

Once you get a digital copy of it you just click the "attach file" button and navigate to the file, etc.

OpenStudy (anonymous):

I hate to sound dumb but do I scan the paper and attach it the computer?

OpenStudy (anonymous):

Do you have a digital camera?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Take a picture and copy the file to your computer, then use the attach button on the website to upload it.

OpenStudy (anonymous):

Well my camera is acting up!!!! I'm so frustrated!!

OpenStudy (anonymous):

Just ask her teacher. They're generally pretty understanding about this sort of thing.

OpenStudy (anonymous):

Here is the part 1of the problem. It says: Draw a model that represents each fraction below. The model should look like the connecting cubes you construct. 1/3 and 2/12

OpenStudy (anonymous):

I really appreciate all of your help. I will E-mail her teacher so she can get a better understanding on this assignment. Thank you again.